# Christoph Bandt

### v1

@@ 3 Automata with two states
Figure 1 showed a topological automaton with two states - the initial state will not be counted. It can be easily modified to describe the decimal numbers with $D=\{0,1,...,9\}$: the edge from $o$ to $right$ will have labels $(0,1),(1,2),(2,3),...,(8,9),$ and the loop at state $right$ gets the label $(9,0).$ Similarly we can describe the number system with respect to any positive integer base.
Let us go back to $D=\{0,1\}$ and look for modifications of Figure 1. Because of properties 3 and 4, the labels cannot be changed, except for one label of a loop, say at $right.$ If we replace $(1,0)$ by $(0,1)$ or $(0,0),$ however, the infinite path from $o$ to $right$ and then traversing the loop will determine the sequences $s,t$ with $s=\overline{0}.$ This contradicts Proposition 3. If we take $(1,1)$ for that label, we would have $t=\overline{1}.$ The conclusion is that for the graph of Figure 1 no other labelling defines a topological automaton.
@@
So we have to change the graph: we replace the loops by arrows to the inverse state, as shown in Figure 2a. The label $(i,j)$ for the arrow from $c$ to $b$ can be chosen, the other must be $(j,i).$ If we take $(0,0)$ or $(1,1),$ we shall again have sequences $s,t$ with $s=\overline{0}$ or $t=\overline{1}.$ The label $(1,0)$ would give $s=\overline{01}.$ The only possible label for the edge $(c,b)$ is $(0,1).$ This automaton describes the number system with base $-2,$ with address map $\varphi(i_{1}i_{2}...)=\sum_{k=1}^{\infty}i_{k}\cdot(-2)^{-k}$ and $X=[-\frac{2}{3},\frac{1}{3}].$ The ‘identification formula’ is $0\overline{01}\sim 1\overline{10}$ while it was $0\overline{1}\sim 1\overline{0}$ for binary numbers.
The third graph has two self-inverse states $b,c.$ Without loss of generality there is a double arrow from $o$ to $b$ which must have labels $(0,1)$ and $(1,0),$ because of properties 2-4. There must be an arrow from $b$ to $c.$ Arrows between self-inverse states can only have labels of the form $(i,i).$ Let us take the label $(1,1)$ for the edge $(b,c).$ By property 1 there is an arrow starting in $c.$ If it leads back to $b,$ we obtain a contradiction with periodic $s.$ So this edge must be a loop, with label $(0,0)$ to avoid periodicity. See Figure 2b. The identification formula is $01\overline{0}\sim 11\overline{0}.$ This automaton represents the symbolic dynamics of the tent map $T:[0,1]\to[0,1],\ T(x)=2x$ for $x\leq\frac{1}{2}$ and $T(x)=2(1-x)$ for $x\geq\frac{1}{2}.$ We have $\varphi^{-1}(x)=j_{0}j_{1}j_{2}...$ with $j_{k}=0$ if $T^{k}(x)\leq\frac{1}{2}$ and $j_{k}=1$ if $T^{k}(x)\geq\frac{1}{2}.$ The critical point $x=\frac{1}{2}$ has two addresses. The tent map is conjugate to the quadratic function $g(x)=x^{2}-2$ on its Julia set $[-2,2]$ so $X$ could be called a Julia set in this case. An equivalent automaton is obtained by exchanging $(0,0)$ and $(1,1)$ for which we have to put the tent map upside down. We thus extended an old result of Hata [22, p. 399] and Bandt and Keller [5, Proposition 4] to automata-generated spaces.
Proposition 4
There are three different topological automata with two states (not counting $o$) and a two-digit alphabet. They describe binary numbers, the number system with base $-2$ and the symbolic dynamics of the tent map. In all cases, the generated space is an interval. normal-$\Box$
@@
Now let us take alphabets with more than two digits. When we add a digit 2 in Figure 1 without adding any labels with 2, except for the mandatory loop $(2,2)$ at $o,$ we obtain a space $X$ where $X_{2}$ is disconnected from $X_{0}\cup X_{1},$ and $X_{02}$ is disconnected from $X_{00}\cup X_{01}$ etc. One metric realization is sketched in Figure 3. The space consists of countably many intervals and a continuum of isolated points since it has Hausdorff dimension greater than 1. In this way, any connected space can be transformed into an archipelago by adding one more digit. For this reason, we focus our study on connected spaces. Disconnected spaces described by two-state automata and their Lipschitz equivalence were studied by Zhu and Yang [51].
@@
Proposition 5
Let $G$ be a topological automaton with $m$ digits and two states which are inverse to each other. Suppose that $G$ completely describes the equivalence of addresses and that $X$ is connected. Then $X$ is an interval, and $G$ defines the double addresses of a number system, either with base $m$ or with $-m.$
Proof. We consider only the case of Figure 1. The case of Figure 2a is similar. Consider an undirected graph $H$ with vertex set $D$ and with edges ${i,j}$ for all labels $(i,j)$ or $(j,i)$ of the edge from $o$ to right. A label $(i,i)$ is not possible since then the state right would be self-inverse. Since $X$ is connected, the graph $H$ is connected and thus has at least $m-1$ edges. Since $G$ is complete, the first and second coordinates of the labels must be different, as noted above. That implies that $H$ has exactly $m-1$ edges which form an undirected path. By renaming the digits we can assume that the labels are $(0,1),(1,2),...,(m-2,m-1).$ Then the loop at state right can have only label $(m-1,0)$ because of the Proposition 3. This gives the number system with base $m,$ with identifications $k\overline{m-1}\sim(k+1)\overline{0}$ for $k=0,...,m-2.$Box
@@
Example 2 (An exotic space)
Figure 5 shows an automaton with two self-inverse states beside $o$ and symbols $0,1,2$ All multiple addresses come as triplets. When we cancel the label $(2,2)$ at $c,$ then $X$ will be the dendroid at the top right in Figure 4. The label $(2,2)$ makes the space more complicated since $X_{0}\cap X_{1}\cap X_{2}$ is a Cantor set, given by all triple addresses $01s\sim 11s\sim 21s$ with $s\in\{0,2\}^{\infty}.$ Actually, $X$ is not homeomorphic to a subset of the plane, as we shall prove in Proposition 8. Figure 5 shows graph approximations of $X$ where the vertex set is $D^{n}$ and two vertices $u,v$ are connected by an edge if $(u,v)$ is accepted by the automaton. Such approximations will be discussed in Section 7.
Figure 5: A complete automaton with 2 self-inverse states and 3 digits. Top right: Without label $(2,2)$ at $c,$ a dendrite $X$ is generated, sketched here on level 5. Bottom: The full space $X$ cannot be embedded into the plane. Graph approximations on level 3 and 4 are shown.
@@ 4 Topological self-similarity of automata-generated spaces
This implies that there are homeomorphisms $h_{i}:X\to X_{i}$ for $i\in D$ such that $h_{i}\cdot\varphi=\varphi\cdot\tau_{i}$ where $\tau_{i}:S\to S$ is defined by $\tau_{i}(s_{1}s_{2}...)=is_{1}s_{2}...\,.$ They fulfil the self-similarity equation
$X=h_{0}(X)\cup...\cup h_{m-1}(X)\ .$ (2)
@@
Proof. The first assertion concerns the infinite paths through loops at $o.$ Now fix $i\in D.$ Any path of edges starting in $o$ can be augmented by putting the loop $(i,i)$ in front of the path. On the other hand, if the path starts with the label $(i,i),$ then this must be the loop at $o$ by property 2, and we get another path by cancelling this loop. Thus each equivalence class of addresses is mapped by $\tau_{i}$ onto another equivalence class. Now $h_{i}$ is just this mapping $\tau_{i}$ but acting on equivalence classes instead of single addresses. It is the morphism of the quotient space $X$ induced by $\tau_{i},$ formally written as $h_{i}(x)=\varphi^{-1}\tau_{i}\varphi(x)$ for $x\in X.$
@@
The topological self-similarity is obvious for the symbolic space $S$ and can be written as equation $S=\tau_{0}(S)\cup...\cup\tau_{m-1}(S)$ which describes a disjoint union. For the quotient space $X$ this implies (2) where the intersections of the $X_{i}=h_{i}(X)$ are usually not empty. They are subject to the rules of the automaton, however. $\Box$
@@
Note that self-similarity of number systems is just a matter of convenience. We treat all cylinders $S_{u}$ of $S$ in the same way, identifying associated pairs of addresses and performing the same operations. It would be much more work to have specific rules for each cylinder. Here we have assumed property 4 which implies (2). However, it turns out that some weaker type of self-similarity directly follows from the generation of $X$ by an automaton.
Theorem 1 (Self-similarity of automata-generated spaces)
@@
Suppose that in Definition 1 we replace the property 4 by the necessary requirement that $G$ accepts all pairs $(u,u)$ of equal words. Then the generated space $X$ and its pieces satisfy some graph-directed topological self-similarity equations.
@@
Proof. Let $V_{0}$ be the set of all states $c$ for which there is a path of edges starting in $o$ with labels $(i_{1},i_{1}),(i_{2},i_{2}),...,(i_{n},i_{n}).$ By property 2, this path will be the only way to accept the pair $(u,u)$ for $u=i_{1}...i_{n}.$ Since we assume also $(ui,ui)$ to be accepted for $i\in D,$ each such state $c$ must have outgoing edges with label $(i,i)$ for each digit $i\in D.$ Since we do not require loops, this means that each $c\in V_{0}$ can be taken as initial state of the automaton and will then generate a topological space $X^{c}$ and an address map $\varphi_{c}:S\to X^{c}.$ The edges of $V_{0}$ will determine how these spaces and maps depend on each other.
@@
Figure 6: Graph-directed topological self-similarity for an automaton without property 4. When we start in state $c,$ we obtain an interval. When we start in $o,$ the space $X$ consists of a sequence of intervals $X_{1},X_{01},X_{001},...$ and a limit point with address $\overline{0}.$ The spiral arrangement of intervals was chosen for good visibility.
@@
Example. Figure 6 shows a very simple example with $V_{0}=\{o,c\}.$ The outgoing edges from $c$ are as in Figure 1 so that with initial state $c$ we get the interval $X^{c}=[0,1]$ and $\varphi_{c}:S\to X^{c}$ identifies the addresses of binary numbers. Proposition 6 holds because of the two loops at $c.$ So we have two homeomorphisms $h^{c}_{i}:X^{c}\to X^{c}_{i}.$ Equation (2) becomes $X^{c}=h^{c}_{0}(X^{c})\cup h^{c}_{1}(X^{c}).$
@@
At the initial state $o$ we have no outgoing edges labelled $(i,j)$ with $i\not=j$ so that our Definition 1 with loop $(1,1)$ at $o$ would yield no identifications and $\varphi:S\to X=X^{id}$ would be the identity map. Indeed $X_{0}$ and $X_{1}$ are disjoint. However, the edge with label $(1,1)$ from $o$ goes to $c.$ It says that there are identifications in $X_{1}$ exactly as in $X^{c}$. Thus $X_{1}$ is an interval, and so is $X_{0^{k}1}$ for $k=1,2,...,$ as seen in Figure 6. Actually, $X$ has almost the same identifications as $X^{c}.$ Only the links $0^{k}0\overline{1}\sim 0^{k}1\overline{0}$ between successive intervals are missing for $k=0,1,2,...\,.$
@@
Since identifications in $X^{c}$ are the same as in the piece $X_{1},$ there is a natural homeomorphism $h_{1}:X^{c}\to X_{1}\subseteq X$ which can be written as $h_{1}(y)=\varphi\tau_{i}\varphi_{c}^{-1}(y)$ for $y\in X^{c}.$ Note that $h_{1}$ corresponds to the edge from $o$ to $c$ but it maps into the opposite direction. Moreover, identifications in the piece $X_{0}$ are the same as in $X,$ so as in Proposition 6 we have the homeomorphism $h_{0}$ from $X$ onto $X_{0}.$ Summarizing we get a system of set equations
@@
Proof continued. In the general case there is one quotient space $X^{c}$ for each possible initial state $c\in V_{0}.$ And we said that for each $c$ and each digit $i\in D$ we have a unique edge $e(c,i)$ starting in $c$ with label $(i,i),$ with a uniquely determined endpoint $d(c,i)\in V_{0}.$ Thus starting in $c$ with digit $i,$ we will make the same identifications as in $X^{d(c,i)}.$ Thus there is a homeomorphism $h_{i}^{c}:X^{d(c,i)}\to X_{i}^{c}$ onto the piece $X_{i}^{c}$ of $X^{c}.$ This yields the equations
@@
This system of equations expresses the graph-directed self-similarity as introduced by Mauldin and Williams [35] and other authors [20, 3, 11, 15] for contractive similitudes $h_{i}^{c}$ in metric spaces. Contractive maps are needed to prove that there is a solution consisting of compact sets $X^{c},c\in V_{0}.$ In our case the solution is constructed by the automaton. So we can consider the more general case of homeomorphisms. $\Box$
@@
This raises mathematical questions which we only mention. What is the appropriate isomorphism concept for automata-generated spaces? Homeomorphy is to wide, Lipschitz equivalence [33, 40, 50, 51] seems to narrow. What about quasiconformal maps? Or shall we better speak about spaces with a graded block structure?

### v2

@@ 3 Automata with two states
Figure 1 showed a topological automaton with two states - the initial state will not be counted. It can be easily modified to describe the decimal numbers with $D=\{0,1,\dots,9\}$: the edge from $o$ to $right$ will have labels $(0,1),(1,2),(2,3),\dots,(8,9)$, and the loop at state $right$ gets the label $(9,0)$. Similarly we can describe the number system with respect to any positive integer base.
Let us go back to $D=\{0,1\}$ and look for modifications of Figure 1. Because of properties 3 and 4, the labels cannot be changed, except for one label of a loop, say at $right$. If we replace $(1,0)$ by $(0,1)$ or $(0,0)$, however, the infinite path from $o$ to $right$ and then traversing the loop will determine the sequences $s,t$ with $s=\overline{0}$. This contradicts Proposition 2.7. If we take $(1,1)$ for that label, we would have $t=\overline{1}$. The conclusion is that for the graph of Figure 1 no other labeling defines a topological automaton.
@@
So we have to change the graph: we replace the loops by arrows to the inverse state, as shown in Figure 2a. The label $(i,j)$ for the arrow from $c$ to $b$ can be chosen, the other must be $(j,i)$. If we take $(0,0)$ or $(1,1)$, we shall again have sequences $s,t$ with $s=\overline{0}$ or $t=\overline{1}$. The label $(1,0)$ would give $s=\overline{01}$. The only possible label for the edge $(c,b)$ is $(0,1)$. This automaton describes the number system with base $-2$, with address map $\varphi(i_{1}i_{2}\dots)=\sum_{k=1}^{\infty}i_{k}\cdot(-2)^{-k}$ and $X=[-\frac{2}{3},\frac{1}{3}]$. The ‘identification formula’ is $0\overline{01}\sim 1\overline{10}$ while it was $0\overline{1}\sim 1\overline{0}$ for binary numbers.
The third graph has two self-inverse states $b,c$. Without loss of generality there is a double arrow from $o$ to $b$ which must have labels $(0,1)$ and $(1,0)$, because of properties 2-4. There must be an arrow from $b$ to $c$. Arrows between self-inverse states can only have labels of the form $(i,i)$. Let us take the label $(1,1)$ for the edge $(b,c)$. By property 1 there is an arrow starting in $c$. If it leads back to $b$, we obtain a contradiction with periodic $s$. So this edge must be a loop, with label $(0,0)$ to avoid periodicity. See Figure 2b. The identification formula is $01\overline{0}\sim 11\overline{0}$. This automaton represents the symbolic dynamics of the tent map $T:[0,1]\to[0,1],\ T(x)=2x$ for $x\leq\frac{1}{2}$ and $T(x)=2(1-x)$ for $x\geq\frac{1}{2}$. We have $\varphi^{-1}(x)=j_{0}j_{1}j_{2}\dots$ with $j_{k}=0$ if $T^{k}(x)\leq\frac{1}{2}$ and $j_{k}=1$ if $T^{k}(x)\geq\frac{1}{2}$. The critical point $x=\frac{1}{2}$ has two addresses. The tent map is conjugate to the quadratic function $g(x)=x^{2}-2$ on its Julia set $[-2,2]$ so $X$ could be called a Julia set in this case. An equivalent automaton is obtained by exchanging $(0,0)$ and $(1,1)$ for which we have to put the tent map upside down. We thus extended an old result of Hata [23, p. 399] and Bandt and Keller [7, Proposition 4] to automata-generated spaces.
Proposition 3.1.
There are three different topological automata with two states (not counting $o$) and a two-digit alphabet. They describe binary numbers, the number system with base $-2$ and the symbolic dynamics of the tent map. In all cases, the generated space is an interval. $\Box$
@@
Now let us take alphabets with more than two digits. When we add a digit 2 in Figure 1 without adding any labels with 2, except for the mandatory loop $(2,2)$ at $o$, we obtain a space $X$ where $X_{2}$ is disconnected from $X_{0}\cup X_{1}$, and $X_{02}$ is disconnected from $X_{00}\cup X_{01}$ etc. One metric realization is sketched in Figure 3. The space consists of countably many intervals and a continuum of isolated points since it has a Hausdorff dimension greater than 1. In this way, any connected space can be transformed into an archipelago by adding one more digit. For this reason, we focus our study on connected spaces. Disconnected spaces described by two-state automata and their Lipschitz equivalence were studied by Zhu and Yang [53].
@@
Proposition 3.2.
Let $G$ be a topological automaton with $m$ digits and two states that are inverse to each other. Suppose that $G$ completely describes the equivalence of addresses and that $X$ is connected. Then $X$ is an interval, and $G$ defines the double addresses of a number system, either with base $m$ or with $-m$.
Proof 3.3.
We consider only the case of Figure 1. The case of Figure 2a is similar. Consider an undirected graph $H$ with vertex set $D$ and with edges ${i,j}$ for all labels $(i,j)$ or $(j,i)$ of the edge from $o$ to right. A label $(i,i)$ is not possible since then the state right would be self-inverse. Since $X$ is connected, the graph $H$ is connected and thus has at least $m-1$ edges. Since $G$ is complete, the first and second coordinates of the labels must be different, as noted above. That implies that $H$ has exactly $m-1$ edges which form an undirected path. By renaming the digits we can assume that the labels are $(0,1),(1,2),\dots,(m-2,m-1)$. Then the loop at state right can have only label $(m-1,0)$ because of the Proposition 2.7. This gives the number system with base $m$, with identifications $k\overline{m-1}\sim(k+1)\overline{0}$ for $k=0,\dots,m-2$.
@@
Example 3.4 (An exotic space).
Figure 5 shows an automaton with two self-inverse states beside $o$ and symbols $0,1,2$. All multiple addresses come as triplets. When we cancel the label $(2,2)$ at $c$, then $X$ will be the dendroid at the top right in Figure 4. The label $(2,2)$ makes the space more complicated since $X_{0}\cap X_{1}\cap X_{2}$ is a Cantor set, given by all triple addresses $01s\sim 11s\sim 21s$ with $s\in\{0,2\}^{\infty}$. Actually, $X$ is not homeomorphic to a subset of the plane, as we shall prove in Proposition 8.1. Figure 5 shows graph approximations of $X$ where the vertex set is $D^{n}$ and two vertices $u,v$ are connected by an edge if $(u,v)$ is accepted by the automaton. Such approximations will be discussed in Section 7.
Figure 5: A complete automaton with 2 self-inverse states and 3 digits. Top right: Without label $(2,2)$ at $c$, a dendrite $X$ is generated, sketched here on level 5. Bottom: The full space $X$ cannot be embedded into the plane. Graph approximations on levels 3 and 4 are shown.
@@ 4 Topological self-similarity of automata-generated spaces
This implies that there are homeomorphisms $h_{i}:X\to X_{i}$ such that $h_{i}\cdot\varphi=\varphi\cdot\tau_{i}$ where $\tau_{i}:S\to S$ is defined by $\tau_{i}(s_{1}s_{2}\dots)=is_{1}s_{2}\dots\,$ for all $i\in D$. They fulfil the self-similarity equation
$X=h_{0}(X)\cup\dots\cup h_{m-1}(X).$ (2)
Proof 4.2.
@@
The first assertion concerns the infinite paths through loops at $o$. Now fix $i\in D$. Any path of edges starting in $o$ can be augmented by putting the loop $(i,i)$ in front of the path. On the other hand, if the path starts with the label $(i,i)$, then this must be the loop at $o$ by property 2, and we get another path by cancelling this loop. Thus each equivalence class of addresses is mapped by $\tau_{i}$ onto another equivalence class. Now $h_{i}$ is just this mapping $\tau_{i}$ but acting on equivalence classes instead of single addresses. It is the morphism of the quotient space $X$ induced by $\tau_{i}$, formally written as $h_{i}(x)=\varphi\tau_{i}\varphi^{-1}(x)$ for $x\in X$.
@@
The topological self-similarity is obvious for the symbolic space $S$ and can be written as equation $S=\tau_{0}(S)\cup\dots\cup\tau_{m-1}(S)$ which describes a disjoint union. For the quotient space $X$ this implies (2) where the intersections of the $X_{i}=h_{i}(X)$ are usually not empty. They are subject to the rules of the automaton, however.
@@
Note that the self-similarity of number systems is just a matter of convenience. We treat all cylinders $S_{u}$ of $S$ in the same way, identifying associated pairs of addresses and performing the same operations. It would be much more work to have specific rules for each cylinder. Here we have assumed property 4 which implies (2). However, it turns out that some weaker type of self-similarity directly follows from the generation of $X$ by an automaton.
Theorem 4.3 (Self-similarity of automata-generated spaces).
@@
Suppose that in Definition 2.4 we replace the property 4 by the requirement that $G$ accepts all pairs $(u,u)$ of equal words. Then the generated space $X$ and its pieces satisfy some graph-directed topological self-similarity equations.
Proof 4.4.
@@
Let $V_{0}$ be the set of all states $c$ for which there is a path of edges starting in $o$ with labels $(i_{1},i_{1}),(i_{2},i_{2}),\dots,(i_{n},i_{n})$. By property 2, this path will be the only way to accept the pair $(u,u)$ for $u=i_{1}\dots i_{n}$. Since we assume also $(ui,ui)$ to be accepted for $i\in D$, each such state $c$ must have outgoing edges with label $(i,i)$ for each digit $i\in D$. Since we do not require loops, this means that each $c\in V_{0}$ can be taken as the initial state of the automaton and will then generate a topological space $X^{c}$ and an address map $\varphi_{c}:S\to X^{c}$. The edges of $V_{0}$ will determine how these spaces and maps depend on each other.
@@
Figure 6: Graph-directed topological self-similarity for an automaton without property 4. When we start in state $c$, we obtain an interval. When we start in $o$, the space $X$ consists of a sequence of intervals $X_{1},X_{01},X_{001},\dots$ and a limit point with address $\overline{0}$. The spiral arrangement of intervals was chosen for good visibility.
Example 4.5.
@@
Figure 6 shows a very simple example with $V_{0}=\{o,c\}$. The outgoing edges from $c$ are as in Figure 1 so that with initial state $c$ we get the interval $X^{c}=[0,1]$ and $\varphi_{c}:S\to X^{c}$ identifies the addresses of binary numbers. Proposition 4.1 holds because of the two loops at $c$. So we have two homeomorphisms $h^{c}_{i}:X^{c}\to X^{c}_{i}$. Equation (2) becomes $X^{c}=h^{c}_{0}(X^{c})\cup h^{c}_{1}(X^{c})$.
@@
At the initial state $o$ we have no outgoing edges labelled $(i,j)$ with $i\not=j$ so that our Definition 2.4 with loop $(1,1)$ at $o$ would yield no identifications and $\varphi:S\to X=X^{o}$ would be the identity map. Indeed $X_{0}$ and $X_{1}$ are disjoint. However, the edge with label $(1,1)$ from $o$ goes to $c$. It says that there are identifications in $X_{1}$ exactly as in $X^{c}$. Thus $X_{1}$ is an interval, and so is $X_{0^{k}1}$ for $k=1,2,\dots$, as seen in Figure 6. Actually, $X$ has almost the same identifications as $X^{c}$. Only the links $0^{k}0\overline{1}\sim 0^{k}1\overline{0}$ between successive intervals are missing for $k=0,1,2,\dots\,$.
@@
Since identifications in $X^{c}$ are the same as in the piece $X_{1}$, there is a natural homeomorphism $h_{1}:X^{c}\to X_{1}\subseteq X$ which can be written as $h_{1}(y)=\varphi\tau_{i}\varphi_{c}^{-1}(y)$ for $y\in X^{c}$. Note that $h_{1}$ corresponds to the edge from $o$ to $c$ and maps into the opposite direction. Moreover, identifications in the piece $X_{0}$ are the same as in $X$, so as in Proposition 4.1 we have the homeomorphism $h_{0}$ from $X$ onto $X_{0}$. Summarizing we get a system of set equations
@@
In the general case there is one quotient space $X^{c}$ for each possible initial state $c\in V_{0}$. And we said that for each $c$ and each digit $i\in D$ we have a unique edge $e(c,i)$ starting in $c$ with label $(i,i)$, with a uniquely determined endpoint $d(c,i)\in V_{0}$. Thus starting in $c$ with digit $i$, we will make the same identifications as in $X^{d(c,i)}$. Thus there is a homeomorphism $h_{i}^{c}:X^{d(c,i)}\to X_{i}^{c}$ onto the piece $X_{i}^{c}$ of $X^{c}$. This yields the equations
@@
This system of equations expresses the graph-directed self-similarity as introduced by Mauldin and Williams [36] and other authors [4, 12, 16, 21] for contractive similitudes $h_{i}^{c}$ in metric spaces. Contractive maps are needed to prove that there is a solution consisting of compact sets $X^{c},c\in V_{0}$. In our case, the solution is constructed by the automaton. So we can consider the more general case of homeomorphisms.
@@
This raises mathematical questions which we only mention. What is the appropriate isomorphism concept for automata-generated spaces? Homeomorphy is to wide, Lipschitz equivalence [34, 42, 52, 53] seems too narrow. What about quasiconformal maps? Or shall we better speak about spaces with a graded block structure?